### v1

# Objective-Free Local Learning and Emergent Language Structure in Thinking Machines

Abstract.
We introduce a neuro-symbolic framework for generative language modeling, based on local event-drive emergent learning. The framework is centered around hierarchical Hopfield memory chains, which act as a compositional short-term memory & dynamic tokenizer (def. retokenizer). Rather than relying on predefined tokens or supervised training, the model builds structure from scratch, learning to retokenize symbol sequences into multi-scale, compositional representations. It does this by hierarchically constructing projection tensors, which glue together symbol sequences into multi-scale tokens. This imbues the model with an inherit redundancy (i.e an emergent gauge structure), and allows the model to compress local feature co-activations into long-range dependencies.
Curiously, we find that the retokenizer can filter natural language patterns from noise, generating synthetic languages with coherent internal morphology – quantifiably the same as human language. Language is learned in a local (Hebbian) fashion, where model constraints dictate allowed emergent structure, and new information is retained in alignment with this structure.
The lack of global objective allows for a type of plasticity not seen in other language models, including allowing it to generalize beyond its initial inference class (and can do so without explicit data). We show that briefly exposing a single new neuron to the retokenizer during inference, binds the distributed multi-scale token features into a symbolic embedding neuron. These embedding neurons behave as an emergent long-term memory, and introduces a key-value memory mechanism that enables compositional inference and generalization.
This architecture provides a methodological foundation for studying how symbolic structure can emerge from local neural learning. It offers a new pathway for building scalable, interpretable neuro-symbolic systems — where tokens, grammar, and reasoning arise as compressed memory traces within a fully unsupervised Hopfield hierarchy. This approach advances the development of scalable, neuromorphic architectures for generative language models.
Exotic AI, Hebbian, Hopfield Networks, Physics-inspired Neural Nets, Interpretable Models, Hierarchical Compression, Language Emergence, Neuro-symbolic, Neuromorphic
1. Introduction
Objective-free local learning is the paradigm by which much of the observed structure and organization in the universe arises. In this setting, learning is governed solely by local rules—no global objective functions, no central coordination. Such systems evolve asynchronously, with higher-order and longer-range correlations emerging as incidental consequences of local interactions unfolding over time.
This paradigm includes many natural and artificial systems: human cognition, biochemical self-organization, and even simple cellular automata such as Conway’s Game of Life. Crucially, learning can emerge from random initial conditions without prestructured input. This implies that structure can be learned in a truly data-free regime, where the learning system constructs its own vocabulary and concepts directly from noise.
In a companion paper (Eugenio, 2025), we derived an effective human language model based on this principle: a hierarchical-chain Hopfield network trained via local Hebbian updates. There, we showed how unsupervised hierarchical feature construction can filter linguistic patterns from noise, giving rise to morphologically coherent vocabularies without supervision or data.
Here, we lay out the broader methodological framework inspired by that model. We describe how local learning events can organize into a recursive projection hierarchy, how memory emerges through replay, and how the resulting system supports key capacities of generative language modeling—tokenization, compositional structure, fast autoregressive inference, and compression—without requiring a global objective, backpropagation, nor explicit data.
This framework also bridges two traditionally distinct approaches to language modeling: correlation-based neural systems (Vaswani et al., 2017; Kaplan et al., 2020; Ramsauer et al., 2021; Kozachkov et al., 2023; Hopfield, 1982; Schultz Kisby et al., 2024) and symbolic compositional reasoning (Lake et al., 2017; d’Avila Garcez and Lamb, 2020; Colelough and Regli, 2025; Lamb et al., 2021; West et al., 2024; Dong et al., 2019). At the lower levels, the model constructs a directed acyclic graph (DAG) of valid $n$-gram combinations, gradually building a recursively compositional vocabulary through a process we call retokenization. Each layer of the hierarchy learns local statistical correlations–akin to identifying relevant interactions in a renormalization group flow (Wilson and Kogut, 1974; Shankar, 1994, 2017)–and projects them into higher-level tokens. In this way, the system builds structured representations across scales, where short-range associations give rise to long-range compositional units.
This sub-symbolic mechanism allows the model to infer token boundaries without explicit markers (e.g space or stop tokens), defining a notion of string acceptance grounded in learned statistical compatibility. The projector maps encode a hierarchy of memory that approximates automaton-like behavior [^1] Formally, regular languages are recognized by finite-state automata that lack explicit memory for counting or nesting (Hopcroft et al., 2001). The early stages of our model resemble finite-state structures through local $n$-gram interactions. However, continual learning causes the model to generalize beyond its initial inference class, and may implicitly support more complex behaviors beyond finite-state expressivity (Pérez et al., 2019; Qiu et al., 2025; Rendell, 2011; Pérez et al., 2021; Rendell, 2014; Hahn, 2020; Weiss et al., 2018). A rigorous classification of the exact formal grammar class supported by this model remains a promising direction for future work.: a sequence is tokenized only if it conforms to the model’s internal structure, and next-token predictions follow trajectories constrained by these memory-induced transitions. Thus, inference behaves like a grammar operating over a learned token space, where the distinction between automaton, grammar, & memory storage/retrieval is blurred. Rather than explicitly storing rules, the system implements them implicitly through its architecture, enabling the replay and recombination of stored sequences without relying on explicit symbolic guidance.
Critically, the model acquires long-term memory not through fixed architecture but through a dynamic inference-driven learning process we call replay. This mechanism enables the system to regenerate and reorganize previously learned structures, binding them into higher-level embedding neurons. These embeddings act as a sparse key-value memory (Gershman et al., 2025; Ba et al., 2016; Tyulmankov et al., 2021), supporting pattern completion and selective recall. Once replay is used to compress and store learned patterns, the lower-layer short-term memory can be forgotten or reused — a form of memory indirection that enables continual learning and concept recombination. In this way, the model achieves Turing-like memory dynamics (Graves et al., 2014, 2016): hierarchical information can persist even as local memory is overwritten, supporting compositional generalization and long-range inference.
A central strength of this framework lies in its interpretability. Each step in the learning process — from Hebbian updates to hierarchical token composition — is locally defined and analytically tractable. The structure and evolution of the model can be explicitly visualized through small toy examples, enabling direct inspection of how words are learned, compressed, and replayed. Unlike black-box models, the memory tensors and projector maps in each layer can be examined and manipulated, allowing one to trace the flow of information across layers and interpret model decisions in terms of compositional features. This makes the framework not only an effective generative model but also a platform for theoretical analysis and conceptual clarity — a rare intersection of symbolic transparency and neural dynamics.
Because of this, we have organized the first half of the paper (Sec 2-6) as both the model introduction & practical tutorial. Sec 2-3 discuss short-term memory (STM). Sec 4-5 discusses long-term memory (LTM) & their emergent key-value mechanisms. In Sec 7, we demonstrate how compositional STM can be creative. We touch on the deeper relationship between inductive biases and natural language structure in Sec 7.1. We end with an extended discussion.
2. Model Summary: Local Events and Hierarchical Retokenization
2.1. Local Learning Events
The foundational object of our framework is a localized event, defined as a pair
$\mathcal{E}^{(n)}=\big{(}\tilde{v}^{(n-1)}_{\mu_{n-1}},\ v_{k}\big{)},$
where $v_{k}$ is the free token to be predicted, and $\tilde{v}^{(n-1)}_{\mu_{n-1}}$ is the learned projection of the $n{-}1$ trailing context tokens. Each such event represents a local time-step in the dynamics of the model, and triggers a Hebbian memory update:
(1) $g^{(n)}_{\mu_{n-1},k}\leftarrow g^{(n)}_{\mu_{n-1},k}+\tilde{v}^{(n-1)}_{\mu_{n-1}}v_{k}.$
This update defines a short-term memory synapse, storing pairwise correlations between projected context features and the next token. Learning is entirely local and unsupervised: there is no global loss function, and no backpropagation.
2.2. Hierarchical Retokenization
To organize these events across scales, the model constructs a recursive sequence of projection maps $P_{n}$ that define higher-order features. Each projection promotes a pair $(\tilde{v}^{(n-1)},v)$ to a new token $\tilde{v}^{(n)}$, if it satisfies the condition:
(2) $\tilde{v}^{(n)}_{\mu_{n}}=\sum_{\mu_{n-1},k}P_{n}^{\mu_{n},\mu_{n-1},k}\tilde{v}^{(n-1)}_{\mu_{n-1}}v_{k},\qquad P_{n}^{\mu_{n},\mu_{n-1},k}\in\{0,1\}.$
These retokenized features form the model’s internal vocabulary. Importantly, the projection maps are constrained by a smoothness condition: a new $n$-gram token is allowed only if all its subcomponents are already valid $(n{-}1)$-grams. This ensures that every token in the hierarchy is recursively decomposable — i.e retokenizable.
2.3. Hierarchical Energy and Inference
The total model energy over all layers is:
(3) $H=\sum_{n\geq 2}\sum_{\mu_{n-1},k}g^{(n)}_{\mu_{n-1},k}\,\tilde{v}^{(n-1)}_{\mu_{n-1}}\,v_{k}.$
Inference corresponds to computing the energy gradient with respect to the next token:
(4) $\frac{\partial H}{\partial v_{k}}=\sum_{n\geq 2}\sum_{\mu_{n-1}}g^{(n)}_{\mu_{n-1},k}\tilde{v}^{(n-1)}_{\mu_{n-1}},$
with update rule $v_{k}\leftarrow f_{k}(\frac{\partial H}{\partial v_{k}})$, where $f$ is a bounded function acting on every element of $v_{k}$. This quantity defines a fast autoregressive prediction mechanism: each token is generated by summing over its correlations with retokenized contexts across multiple scales. The value of the next token is sampled from a probability distribution, e.g $\rho_{k}(v_{k})=e^{v_{k}}/\sum_{j}e^{v_{j}}$. Note, the model can be biased to favor longer-range dependencies by rescaling weights of deeper layers: $g^{(n)}\rightarrow\beta^{n}g^{(n)}$ with $\beta>1$.
3. Retokenization and Emergent $n$-Point Structure
Although each layer of the model stores only local 2-point correlations between a projected context token and a predicted token,
$H_{n}=\sum_{\mu_{n-1},k}g^{(n)}_{\mu_{n-1},k}\,\tilde{v}^{(n-1)}_{\mu_{n-1}}\,v_{k},$
the recursive definition of projected tokens induces an implicit $n$-point interaction. Each projection is defined via Eqn 2 and applying this recursively from the basis layer ($n=1$) yields:
(5) $\tilde{v}^{(n-1)}_{\mu_{n-1}}$ $=\sum_{j_{1},\ldots,j_{n-1}}(P_{n-1}\cdots P_{3}P_{2})^{\mu_{n-1}}_{j_{1}\cdots j_{n-1}}v_{j_{1}}v_{j_{2}}\cdots v_{j_{n-1}},$
(6) $\Rightarrow\quad H_{n}$ $=\sum_{j_{1},\ldots,j_{n}}g^{(n)}_{j_{1}\cdots j_{n}}\,v_{j_{1}}\cdots v_{j_{n}},$
where
$g^{(n)}_{j_{1}\cdots j_{n}}=\sum_{\mu_{n-1}}g^{(n)}_{\mu_{n-1},j_{n}}\cdot(P_{n-1}\cdots P_{2})^{\mu_{n-1}}_{j_{1}\cdots j_{n-1}}.$
Thus, a learned 2-point memory update at level $n$ becomes an effective $n$-point interaction through hierarchical retokenization. This allows the model to encode long-range dependencies in a compressed form using only local unsupervised learning events.
3.1. Retokenization Symmetries and Gauge Equivalence
The same tensor $\tilde{v}^{(n)}$ can be expressed with either left- or right-tokenized orderings of its internal structure. These forms are related via recursive retokenization of the projection tensors. Given a right-tokenized projector $P_{n}^{\mu_{n},\mu_{n-1},k}$, we define the corresponding left-tokenized form recursively as:
(7) $P_{L,n}^{\mu_{n},j_{1},\mu_{n-1}}=\sum_{k,\mu_{n-2}}\left(\sum_{\mu_{n-1}^{\prime}}P_{n}^{\mu_{n},\mu_{n-1}^{\prime},k}\cdot P_{L,n-1}^{\mu_{n-1}^{\prime},j_{1},\mu_{n-2}}\right)\cdot P_{n-1}^{\mu_{n-1},\mu_{n-2},k},$
with base case $P_{L,2}^{\mu_{2},j_{1},j_{2}}:=P_{2}^{\mu_{2},j_{1},j_{2}}$. These transformations preserve the smooth gauge structure of the model, and maps us between different representations of the same smooth tensor, which are distinguished by the order & number of indices. (E.g $g^{(n)}_{\mu_{n}}$, $g^{(n)}_{\mu_{n-1},k}$, & $g^{(n)}_{k,\mu_{n-1}}$ are different representations of $g^{(n)}$.) Where reordering is done via retokenization and not transposition ($g^{(n)}_{\mu_{n-1},k}\neq g^{(n)T}_{k,\mu_{n-1}}$ in general). In practice, this means:
$\sum_{\mu_{n-1},k}P_{n}^{\mu_{n},\mu_{n-1},k}\tilde{v}^{(n-1)}_{\mu_{n-1}}v_{k}=\sum_{k,\mu_{n-1}}P_{L,n}^{\mu_{n},k,\mu_{n-1}}v_{k}\tilde{v}^{(n-1)}_{\mu_{n-1}}.$
This formalism ensures that all token representations and learning updates remain strictly local and compositional, while implicitly capturing longer correlations.
We may understand this redundancy of the hierarchical featureset as manifesting an emergent associativity, such that the tensor product $v_{\bf c}v_{\bf a}v_{\bf t}$ can be recast as $\tilde{v}_{\bf cat}^{(3)}$, $\tilde{v}_{\bf ca}^{(2)}v_{\bf t}$, & $v_{\bf c}\tilde{v}_{\bf at}^{(2)}$ via $P_{n}$’s. We exploit this in the following sections.
3.2. Left-Inference via Retokenization
To perform left-inference—predicting the initial token of a word given trailing context—we retokenize the energy function into a left-gauge form. For each layer $n$, the hierarchical Hopfield model learns 2-point correlations of the form:
(8) $H_{n}=\sum_{\mu_{n-1},k}g^{(n)}_{\mu_{n-1},k}\,\tilde{v}^{(n-1)}_{\mu_{n-1}}\,v_{k},$
where $\tilde{v}^{(n-1)}$ is the projected trailing context, and $v_{k}$ is the final (predicted) token. This is the default right-inference form.
To express $H_{n}$ in a form suitable for left-inference, we apply the left-tokenized projector $P_{L,n}^{\mu_{n},j,\mu_{n-1}}$, defined recursively via:
(9) $\tilde{v}^{(n)}_{\mu_{n}}=\sum_{j,\mu_{n-1}}P_{L,n}^{\mu_{n},j,\mu_{n-1}}\,v_{j}\tilde{v}^{(n-1)}_{\mu_{n-1}}.$
The energy becomes:
(10) $H_{n}=\sum_{\mu_{n}}g^{(n)}_{\mu_{n}}\,\tilde{v}^{(n)}_{\mu_{n}}=\sum_{j,\mu_{n-1},\mu_{n}}g^{(n)}_{\mu_{n}}\,P_{L,n}^{\mu_{n},j,\mu_{n-1}}\,v_{j}\tilde{v}^{(n-1)}_{\mu_{n-1}}.$
To compute the gradient for left-inference, we differentiate with respect to the initial token $v_{j}$:
(11) $\frac{\partial H}{\partial v_{j}}=\sum_{n\geq 2}\sum_{\mu_{n-1},\mu_{n}}g^{(n)}_{\mu_{n}}\,P_{L,n}^{\mu_{n},j,\mu_{n-1}}\,\tilde{v}^{(n-1)}_{\mu_{n-1}}.$
This is structurally symmetric to the right-inference form:
(12) $\frac{\partial H}{\partial v_{k}}=\sum_{n\geq 2}\sum_{\mu_{n-1}}g^{(n)}_{\mu_{n-1},k}\,\tilde{v}^{(n-1)}_{\mu_{n-1}},$
but now uses the left-retokenized projector and reversed token ordering.
3.3. Example: Learning from dogs, cats, and frogs
To illustrate how the hierarchical Hopfield model builds structure from scratch, we walk through the learning process for a minimal training set consisting of three words:
${\bf dogs},\quad{\bf cats},\quad{\bf frog}.$
We assume the model begins with a basis token set of dimension $d=26$, one for each letter of the English alphabet: $v_{j}\in\{v_{\bf a},\dots,v_{\bf z}\}$. The model learns from co-occurrence patterns in the training data, using local Hebbian updates following Eqn 1.
Step 1: Learning Bigrams
From the three words, the following bigrams (2-grams) occur:
${\bf do},\ {\bf og},\ {\bf gs},\quad{\bf ca},\ {\bf at},\ {\bf ts},\quad{\bf fr},\ {\bf ro},\ {\bf og}.$
The model defines a projection tensor $P_{2}^{\mu_{2},j_{1},j_{2}}$, with a separate index $\mu_{2}$ for each learned bigram:
$P_{2}=\sum_{\mu_{2}}P_{2}^{\mu_{2},j_{1},j_{2}}\,\tilde{v}^{(2)}_{\mu_{2}}\,v_{j_{1}}v_{j_{2}},$
where $P_{2}^{\mu_{2},j_{1},j_{2}}=1$ for each bigram above. These define eight new 2-gram tokens:
$\tilde{v}^{(2)}_{\bf do},\ \tilde{v}^{(2)}_{\bf og},\ \tilde{v}^{(2)}_{\bf gs},\ \tilde{v}^{(2)}_{\bf ca},\ \tilde{v}^{(2)}_{\bf at},\ \tilde{v}^{(2)}_{\bf ts},\ \tilde{v}^{(2)}_{\bf fr},\ \tilde{v}^{(2)}_{\bf ro}.$
Step 2: Learning Trigrams
The trigrams (3-grams) in the words are:
${\bf dog},\ {\bf ogs},\quad{\bf cat},\ {\bf ats},\quad{\bf fro},\ {\bf rog}.$
From the smooth bigrams above, the model constructs a new projector:
$P_{3}^{\mu_{3},\mu_{2},j_{3}}=1\quad\text{for learned combinations}\quad\tilde{v}^{(3)}_{\mu_{3}}=\tilde{v}^{(2)}_{\mu_{2}}\,v_{j_{3}}.$
This yields six smooth trigram tokens:
$\tilde{v}^{(3)}_{\bf dog},\ \tilde{v}^{(3)}_{\bf ogs},\ \tilde{v}^{(3)}_{\bf cat},\ \tilde{v}^{(3)}_{\bf ats},\ \tilde{v}^{(3)}_{\bf fro},\ \tilde{v}^{(3)}_{\bf rog}.$
Step 3: Learning 4-grams (Full Words)
Finally, the full words dogs, cats, and frog are learned:
$\tilde{v}^{(3)}_{\bf dog}\,v_{\bf s}\quad\Rightarrow\quad\tilde{v}^{(4)}_{\bf dogs},\qquad\tilde{v}^{(3)}_{\bf cat}\,v_{\bf s}\quad\Rightarrow\quad\tilde{v}^{(4)}_{\bf cats},$
$\tilde{v}^{(3)}_{\bf fro}\,v_{\bf g}\quad\Rightarrow\quad\tilde{v}^{(4)}_{\bf frog}.$
Summary: Smooth Tokenset
After training, the smooth tokenset contains:
  • 8 bigrams: $\tilde{v}^{(2)}_{\bf do},\ \tilde{v}^{(2)}_{\bf og},\ \tilde{v}^{(2)}_{\bf gs},\ \tilde{v}^{(2)}_{\bf ca},\ \tilde{v}^{(2)}_{\bf at},\ \tilde{v}^{(2)}_{\bf ts},\ \tilde{v}^{(2)}_{\bf fr},\ \tilde{v}^{(2)}_{\bf ro}$
  • 6 trigrams: $\tilde{v}^{(3)}_{\bf dog},\ \tilde{v}^{(3)}_{\bf ogs},\ \tilde{v}^{(3)}_{\bf cat},\ \tilde{v}^{(3)}_{\bf ats},\ \tilde{v}^{(3)}_{\bf fro},\ \tilde{v}^{(3)}_{\bf rog}$
  • 3 4-grams: $\tilde{v}^{(4)}_{\bf dogs},\ \tilde{v}^{(4)}_{\bf cats},\ \tilde{v}^{(4)}_{\bf frog}$
Final Projectors
The learned projectors take the form:
(13) $P_{2}^{\mu_{2},j_{1},j_{2}}$ $=1\quad\text{for }(\mu_{2},j_{1}j_{2})\in\scriptstyle{\{({\bf do},{\bf d},{\bf o}),\ ({\bf og},{\bf o},{\bf g}),\ ({\bf gs},{\bf g},{\bf s}),\ ({\bf ca},{\bf c},{\bf a}),}$
$\hskip 85.00012pt\scriptstyle{({\bf at},{\bf a},{\bf t}),\ ({\bf ts},{\bf t},{\bf s}),\ ({\bf fr},{\bf f},{\bf r}),\ ({\bf ro},{\bf r},{\bf o})\}}$
$P_{3}^{\mu_{3},\mu_{2},j_{3}}$ $=1\quad\text{for }(\mu_{3},\mu_{2},j_{3})\in\scriptstyle{\{({\bf dog},{\bf do},{\bf g}),\ ({\bf ogs},{\bf og},{\bf s}),\ ({\bf cat},{\bf ca},{\bf t}),}$
$\scriptstyle{\hskip 59.50012pt({\bf ats},{\bf at},{\bf s}),\ ({\bf fro},{\bf fr},{\bf o}),\ ({\bf rog},{\bf ro},{\bf g})\}}$
$P_{4}^{\mu_{4},\mu_{3},j_{4}}$ $=1\quad\text{for }(\mu_{4},\mu_{3},j_{4})\in\scriptstyle{\{({\bf dogs},{\bf dog},{\bf s}),\ ({\bf cats},{\bf cat},{\bf s}),\ ({\bf frog},{\bf fro},{\bf g})\}}$
Or equivalently written as a superposition of tensor product states (a.k.a an entangled state (Schollwöck, 2011)):
(14) $P_{2}$ $=$ $\tilde{v}_{\bf do}^{(2)}v_{\bf d}v_{\bf o}+\tilde{v}_{\bf og}^{(2)}v_{\bf o}v_{\bf g}+\tilde{v}_{\bf gs}^{(2)}v_{\bf g}v_{\bf s}+\tilde{v}_{\bf ca}^{(2)}v_{\bf c}v_{\bf a}+\tilde{v}_{\bf at}^{(2)}v_{\bf a}v_{\bf t}$
$+$ $\tilde{v}_{\bf ts}^{(2)}v_{\bf t}v_{\bf s}+\tilde{v}_{\bf fr}^{(2)}v_{\bf f}v_{\bf r}+\tilde{v}_{\bf ro}^{(2)}v_{\bf r}v_{\bf o}$
(15) $P_{3}$ $=$ $\tilde{v}_{\bf dog}^{(3)}\tilde{v}_{\bf do}^{(2)}v_{\bf g}+\tilde{v}_{\bf ogs}^{(3)}\tilde{v}_{\bf og}^{(2)}v_{\bf s}+\tilde{v}_{\bf cat}^{(3)}\tilde{v}_{\bf ca}^{(2)}v_{\bf t}+\tilde{v}_{\bf ats}^{(3)}\tilde{v}_{\bf at}^{(2)}v_{\bf s}$
$+$ $\tilde{v}_{\bf fro}^{(3)}\tilde{v}_{\bf fr}^{(2)}v_{\bf o}+\tilde{v}_{\bf rog}^{(3)}\tilde{v}_{\bf ro}^{(2)}v_{\bf g}$
(16) $P_{4}$ $=$ $\tilde{v}_{\bf dogs}^{(4)}\tilde{v}_{\bf dog}^{(3)}v_{\bf s}+\tilde{v}_{\bf cats}^{(4)}\tilde{v}_{\bf cat}^{(3)}v_{\bf s}+\tilde{v}_{\bf frog}^{(4)}\tilde{v}_{\bf fro}^{(3)}v_{\bf g}$
We can understand the product of projector tensors $P_{n}P_{n-1}\cdots P_{2}$ as encoding a bundle of finite state transitions (or feature maps), e.g $v_{\bf r}\rightarrow\tilde{v}^{(2)}_{\bf fr}\rightarrow\tilde{v}^{(3)}_{\bf fro}\rightarrow\tilde{v}^{(4)}_{\bf frog}$, which collectively describe all possible paths toward the terminal nodes of the DAG.
Every token in this hierarchy is retokenizable into only previously learned parts. This guarantees that inference can recursively decompose any learned token, and that the vocabulary is self-consistent and locally compositional. For example, given context frog, the model knows to terminate right-inference because no merger of $\tilde{v}^{(4)}_{\bf frog}$ with another token exists as a feature map – defined by non-vanishing action of $P_{n}P_{n-1}\cdots P_{2}$ (e.g Fig 1). Nevertheless, the model can still infer frogs as the best non-smooth construction due to local interactions $g^{(2)}_{\bf og}$ $g^{(3)}_{\bf ogs}$.
This latter subtlety highlights the hybrid sub-symbolic/automaton nature of the model, where tokenization [^2] This dynamic tokenization is distinct from both tokenization (Schuster and Nakajima, 2012; Gage, 1994) & entropy-based approaches to LLM’s (Li et al., 2025; Zhang et al., 2024; Agrawal et al., 2024; Simonds, 2025). LLM’s assume pretokenized input (Phuong and Hutter, 2022; Gage, 1994; Schuster and Nakajima, 2012). plays a dynamical role during inference to determine boundaries, and therefore imbues the model with a notion of accepted string akin to automata. In energetic terms, the accepted strings live in a low-energy manifold (the smooth tokenset), and therefore are the most-likely configurations to be observed given the context. We may therefore think of the automaton as living in the ground state, with its language being the ground state configurations (Reinhart and las Cuevas, 2022). This approximate behavior is made precise by the introduction of projectors (and therefore a consequence of learning higher-order correlations while requiring locality).
Figure 1. Summarizing the relationship between tokenizability, locality, and hierarchy. Three words from a random language (Eugenio, 2025) (shown in blue): yjjfqi, kltilj, & yjjfgsp.
4. The benefits of thinking about it
The hierarchical short-term memory stores learned words as smooth tokens encoded via a bundle of entangled feature maps ($P_{n}\cdots P_{2}$) and local interactions ($g^{(n)}$). This allows for the model to determine word boundaries (Fig 1), as well as retrieve words from context.
However, this places a number of limitations which limit its scalability. First, it learns all feature maps as entangled, regardless of if they share overlapping graphs – this is highly impractical [^3] Because the feature maps are entangled, the projector size scales as the square of the number of learned features ($n$-grams) (Eugenio, 2025). Because the $P_{n}$ are sparse, a tradeoff can be made to improve memory, however at the cost of losing the representational power of dense matrices. Additionally, memories would have to be stored & accessed from a singular large hierarchy, regardless of whether most of the memory traces are unrelated to the given context. And we noticed the total number of $n$-grams in a real text scales as a power-law (Eugenio, 2025).. Further, word length is finite – stored strings are limited to regular-language-like structure.
The model overcomes these limitations by employing a thinking-like process we call replay relearning. This procedure binds the distributed features of a token sequence (e.g a word object) into a new long-term embedding neuron $a_{\alpha}$, thus disentangling the feature maps into a series of smaller word-specific feature maps. These can later participate in inference independent of the original $g^{(n)}$ (which can be forgotten).
4.1. The Replay Game (Model Thinking)
Replay begins by selecting a random basis token $v_{k}$ as seed. From this seed, the short-term memory network performs left and right inference using only the local energy gradients from $g^{(n)}$. Inference proceeds recursively: the next token $v$ is selected based on the energy gradient
$\Delta v_{k}=\sum_{n,\mu_{n-1}}g^{(n)}_{\mu_{n-1},k}\,\tilde{v}^{(n-1)}_{\mu_{n-1}},$
and sampling $\rho_{k}\big{(}f_{k}(\Delta v_{k})\big{)}$. At each step, retokenization ensures that only smooth (i.e compositional) features are activated.
The process continues until the inferred string reaches a natural boundary — a region where no further smooth retokenization is possible. The result is a spontaneously generated word, such as:
$[v_{\bf c},v_{\bf a},v_{\bf t},v_{\bf s}].$
At some point during this replay cycle — whether mid-replay or post-hoc — we introduce a new auxiliary neuron $a_{\alpha}$ from an external source (conceptually, a pool at infinity). We activate it by construction (i.e we force it to fire).
As the replay proceeds, the features $\tilde{v}^{(n)}$ that fire over time become locally Hebbian-correlated with the firing $a_{\alpha}$. That is, at each replay step, the memory weight is updated by:
$m_{\alpha,\mu_{n}}^{(n)}\leftarrow m_{\alpha,\mu_{n}}^{(n)}+\eta\,a_{\alpha}\tilde{v}_{\mu_{n}}^{(n)},$
where $\eta$ is a learning rate.
Alternatively — and equivalently — one can view the full word replay as producing a batch of active features, which are then bound in a single step to the new $a_{\alpha}$ neuron. It encodes the feature content across all $n$-gram layers of the retokenization hierarchy:
[…]
The result is a key-value memory binding (Tyulmankov et al., 2021; Gershman et al., 2025; Ba et al., 2016): the neuron $a_{\alpha}$ becomes a long-term embedding associated with the full replayed word. Over time, this mechanism constructs a library of long-term memory neurons $\{a_{\alpha}\}$, each associated with a unique word and its compositional features. Once learned, these embeddings enable fast inference via pattern completion: a partial word (e.g. a suffix or prefix) can trigger an embedding, which drives next-token prediction.
Benefits of Replay Relearning
  • Compression: Word features are projected onto low-rank memory vectors.
  • Decoupling: Embeddings allow hierarchical memory $g^{(n)}$ to be forgotten or reused.
  • Continual Learning: New embeddings can be learned without overwriting existing ones.
  • Parallelism: Each $a_{\alpha}$ contributes independently to inference.
  • Emergent Computation: The $a_{\alpha}$ contribution to inference generalizes beyond local $n$-gram statistics.
(a) The dynamic game of “replay”: (1) Randomly activate node in short-term memory (STM). (2) Introduce new neuron until hierarchy converges on stored memory, then remove it. (3) Added neurons emerge as long-term memory (LTM).
(b) LEFT: State-transitions of the STM converge on terminal node. Red shows features activated during replay. RIGHT: Embedding neuron ($a_{1}$) learns an emergent hierarchical key-value memory, which is an attention mechanism wrapped around a word-specific n-gram model. BOTTOM: Replay as autoregression with boundary awareness. Red shows activated trailing context tokens.
Figure 2. Dynamically Coding the Local Plastic Structure
5. Emergent Hierarchical Key-Value Memory (Recognition)
Inference proceeds as a sequence of discrete event updates. At each time step, a new token is predicted based on the trailing context through the structure stored in both short-term ($g^{(n)}$) and long-term ($m^{(n)}$) memory. The full Hamiltonian is therefore
(17) $H=\sum\limits_{n}\sum_{\mu_{n-1},k}g_{\mu_{n-1},k}^{(n)}\tilde{v}_{\mu_{n-1}}^{(n-1)}v_{k}+\sum_{\alpha}a_{\alpha}\Big{(}\sum\limits_{n}\sum_{\mu_{n}}m_{\alpha,\mu_{n}}^{(n)}\tilde{v}_{\mu_{n}}^{(n)}\Big{)}.$
5.1. Update Logic
Each token prediction is based on evaluating local gradients:
(18) $Q\leftarrow f_{Q}\left(\Delta Q\right),\qquad\Delta Q:=\frac{\partial H}{\partial Q},$
for $Q\in\{v_{k},\ a_{\alpha}\}$, where $f_{Q}$ is any bounded activation function.
We choose to update $a_{\alpha}$ neurons first. Their updated values are then used to compute the update for $v_{k}$. The corresponding gradients are:
(19) $\Delta a_{\alpha}$ $=\sum_{n}\sum_{\mu_{n}}m_{\alpha,\mu_{n}}^{(n)}\tilde{v}_{\mu_{n}}^{(n)},$
(20) $\Delta v_{k}$ $=\sum_{n}\sum_{\mu_{n-1}}\left(g^{(n)}_{\mu_{n-1},k}+\sum_{\alpha}a_{\alpha}\,m^{(n)}_{\alpha,\mu_{n-1},k}\right)\tilde{v}^{(n-1)}_{\mu_{n-1}}.$
5.2. Key-Value Memory via Fast Embedding Neurons
If $a_{\alpha}$ neurons respond more rapidly than $v_{k}$, we treat them as reaching a fixed point first (Krotov, 2021):
(21) $a_{\alpha}=f_{\alpha}\left(\sum_{n}\sum_{\mu_{n}}m_{\alpha,\mu_{n}}^{(n)}\tilde{v}_{\mu_{n}}^{(n)}\right).$
Plugging into the equation for $\Delta v_{k}$ yields:
(22) $\Delta v_{k}$ $=\sum_{n}\sum_{\mu_{n-1}}g^{(n)}_{\mu_{n-1},k}\,\tilde{v}^{(n-1)}_{\mu_{n-1}}$
(23) $\quad+\sum_{\alpha}f_{\alpha}\left(\sum_{n^{\prime}}\sum_{\mu_{n^{\prime}}^{\prime}}m_{\alpha,\mu_{n^{\prime}}^{\prime}}^{(n^{\prime})}\tilde{v}_{\mu_{n^{\prime}}^{\prime}}^{(n^{\prime})}\right)\cdot\sum_{n}\sum_{\mu_{n-1}}m_{\alpha,\mu_{n-1},k}^{(n)}\tilde{v}^{(n-1)}_{\mu_{n-1}}.$
Eqn 23 defines a hierarchical key-value memory, where:
  • The key is the set of smooth projected context tokens $\tilde{v}^{(n)}$,
  • The embedding neuron $a_{\alpha}$ is selected by the context via $f_{\alpha}(\cdot)$,
  • The value is the distribution over next tokens $v_{k}$, computed via the activated embedding.
This emergent key-value memory mechanism allows long-term associations to be built from local dynamics alone, forming the basis for structured retrieval and generalization in later inference. Thus the model can generalize beyond local $n$-gram statistics, enabling long-range coherence, partial match inference, and efficient lookup. We refer to this process as recognition [^4] Note that recognition can interfere with tokenization. For instance, the repeated string formed from monk$+$eye, i.e …onkeyemonkeyemon…, uniquely tokenizes back into the two words stated, however, the reader may still recognize monkey or key even though they don’t uniquely tokenize the string, highlighting the competition between fast long-term memory effects with slower tokenizing short-term memory (Eugenio, 2025)..
5.2.1. Key-Value Memory Example
Consider Eqn 23 with $g^{(n)}=0$, s.t inference is driven purely by $a_{\alpha}$. Suppose we have two learned word embeddings, $\alpha\in\{1,2\}$, corresponding to:
$a_{1}\leftrightarrow\text{{\bf cats}},\qquad a_{2}\leftrightarrow\text{{\bf dogs}}.$
Then the update becomes:
$\Delta v_{k}=$
$f_{1}\left(\tilde{v}_{\bf ca}^{(2)}m^{(2)}_{1,{\bf ca}}+\tilde{v}_{\bf at}^{(2)}m^{(2)}_{1,{\bf at}}+\tilde{v}_{\bf cat}^{(3)}m^{(3)}_{1,{\bf cat}}+\tilde{v}_{\bf ats}^{(3)}m^{(3)}_{1,{\bf ats}}+\tilde{v}_{\bf cats}^{(4)}m^{(4)}_{1,{\bf cats}}\right)$
×$\quad\times\left(\tilde{v}_{\bf ca}^{(2)}m^{(3)}_{1,{\bf ca},k}+\tilde{v}_{\bf at}^{(2)}m^{(3)}_{1,{\bf at},k}+\tilde{v}_{\bf cat}^{(3)}m^{(4)}_{1,{\bf cat},k}\right)$
$+f_{2}\left(\tilde{v}_{\bf do}^{(2)}m^{(2)}_{2,{\bf do}}+\tilde{v}_{\bf og}^{(2)}m^{(2)}_{2,{\bf og}}+\tilde{v}_{\bf dog}^{(3)}m^{(3)}_{2,{\bf dog}}+\tilde{v}_{\bf ogs}^{(3)}m^{(3)}_{2,{\bf ogs}}+\tilde{v}_{\bf dogs}^{(4)}m^{(4)}_{2,{\bf dogs}}\right)$
(24) ×$\quad\times\left(\tilde{v}_{\bf do}^{(2)}m^{(3)}_{2,{\bf do},k}+\tilde{v}_{\bf og}^{(2)}m^{(3)}_{2,{\bf og},k}+\tilde{v}_{\bf dog}^{(3)}m^{(4)}_{2,{\bf dog},k}\right).$
Note that generically $m^{(1)}_{\alpha,k}$ (monogram) terms also contribute, but we suppress them in this example to preserve space.
If the context consists of $v_{\bf a},v_{\bf t}$, then $\tilde{v}_{\bf at}^{(2)}=1$ and all other projections vanish. The update reduces to:
$\Delta v_{k}=f_{1}\big{(}m^{(2)}_{1,{\bf at}}\big{)}\cdot m^{(3)}_{1,{\bf at},k}.$
This predicts the next token $v_{k}$ with the strongest overlap with ${\bf s}$, due to the binding $m^{(3)}_{1,{\bf at},k}=m^{(3)}_{1,{\bf ats}}\,\delta_{k,{\bf s}}$.
Note that the update of $a_{\alpha}$ inherits information about morphological structure (the projection maps) from its interaction with the $\tilde{v}^{(n)}$. Thus, smooth features in the context trigger the appropriate word embedding via $f_{\alpha}(\cdot)$, which then drives prediction through projection maps. Thus there is no need to further encode positional relationships, which is handled by the projectors. We discuss how this inheritance allows optimizing for word-specific projection maps in the following section.
5.3. Compression via Word-Specific Projector Decomposition
To illustrate how replay leads to compression, we begin with a concrete example. Suppose the model has learned the following projectors:
Bigrams:
(25) $P_{2}=$ $\tilde{v}^{(2)}_{\bf ca}v_{\bf c}v_{\bf a}+\tilde{v}^{(2)}_{\bf at}v_{\bf a}v_{\bf t}+\tilde{v}^{(2)}_{\bf ts}v_{\bf t}v_{\bf s}$
$+\tilde{v}^{(2)}_{\bf do}v_{\bf d}v_{\bf o}+\tilde{v}^{(2)}_{\bf og}v_{\bf o}v_{\bf g}+\tilde{v}^{(2)}_{\bf gs}v_{\bf g}v_{\bf s}$
Trigrams:
(26) $P_{3}=$ $\tilde{v}^{(3)}_{\bf cat}\tilde{v}^{(2)}_{\bf ca}v_{\bf t}+\tilde{v}^{(3)}_{\bf ats}\tilde{v}^{(2)}_{\bf at}v_{\bf s}$
$+\tilde{v}^{(3)}_{\bf dog}\tilde{v}^{(2)}_{\bf do}v_{\bf g}+\tilde{v}^{(3)}_{\bf ogs}\tilde{v}^{(2)}_{\bf og}v_{\bf s}$
4-grams:
(27) $P_{4}=\tilde{v}^{(4)}_{\bf cats}\tilde{v}^{(3)}_{\bf cat}v_{\bf s}+\tilde{v}^{(4)}_{\bf dogs}\tilde{v}^{(3)}_{\bf dog}v_{\bf s}$
Now consider the replay-imprinted $m$-tensors for $\alpha={\bf cats}$:
(28) $m^{(2)}_{\bf cats}$ $=\tilde{v}^{(2)}_{\bf ca}+\tilde{v}^{(2)}_{\bf at}+\tilde{v}^{(2)}_{\bf ts},$
$m^{(3)}_{\bf cats}$ $=\tilde{v}^{(3)}_{\bf cat}+\tilde{v}^{(3)}_{\bf ats},$
$m^{(4)}_{\bf cats}$ $=\tilde{v}^{(4)}_{\bf cats}.$
We now demonstrate compression by contracting $m^{(3)}_{\bf cats}$ with the entangled projector $P_{3}$:
(29) $m^{(3)}_{\bf cats}P_{3}$ $=(\tilde{v}^{(3)}_{\bf cat}+\tilde{v}^{(3)}_{\bf ats})(\tilde{v}^{(3)}_{\bf cat}\tilde{v}^{(2)}_{\bf ca}v_{\bf t}+\tilde{v}^{(3)}_{\bf ats}\tilde{v}^{(2)}_{\bf at}v_{\bf s}$
$\qquad+\tilde{v}^{(3)}_{\bf dog}\tilde{v}^{(2)}_{\bf do}v_{\bf g}+\tilde{v}^{(3)}_{\bf ogs}\tilde{v}^{(2)}_{\bf og}v_{\bf s})$
$=\tilde{v}^{(2)}_{\bf ca}v_{\bf t}+\tilde{v}^{(2)}_{\bf at}v_{\bf s},$
where orthogonality eliminates the contributions from tokens unrelated to cats. The result is a disentangled, word-specific tensor chain:
(30) $=(\tilde{v}^{(3)}_{\bf cat}+\tilde{v}^{(3)}_{\bf ats})(\tilde{v}^{(3)}_{\bf cat}\tilde{v}^{(2)}_{\bf ca}v_{\bf t}+\tilde{v}^{(3)}_{\bf ats}\tilde{v}^{(2)}_{\bf at}v_{\bf s}):=\hat{m}^{(3)}_{\bf cats}\tilde{P}^{\bf cats}_{3}.$
This procedure can be recursively applied to the full projector chain $P_{3}P_{2}$ to yield a compressed set of projectors $\hat{P}_{n}^{\alpha}$ for each word $\alpha$. These retain only the subspaces relevant to $\alpha$, enabling scalable and disentangled memory representations.
Next, we generalize this process to arbitrary word embeddings using SVD-based decomposition of the $m$-capped projection chains.
5.4. Generalized Compression via Projector Contraction
The concrete example above illustrates how word-specific projector chains can be disentangled by pruning irrelevant branches. We now formalize this procedure using a sequence of algebraic contractions and SVDs (Schollwöck, 2011; Parker et al., 2020; Gourianov et al., 2022), resulting in a compressed representation of the hierarchical projector chain for each word embedding.
We begin with the full replay expression for a given word embedding $a_{\alpha}$:
(31) $\sum_{\mu_{n}}m^{(n)}_{\alpha,\mu_{n}}\,\tilde{v}^{(n)}_{\mu_{n}}=\sum_{\mu_{n},\mu_{n-1},\dots,\mu_{2}}m^{(n)}_{\alpha,\mu_{n}}P_{n}^{\mu_{n},\mu_{n-1},j_{n}}\cdots P_{2}^{\mu_{2},j_{1},j_{2}}\,v_{j_{1}}v_{j_{2}}\cdots v_{j_{n}}.$
This expression can be interpreted as a long chain of projection maps contracted with the memory vector $m^{(n)}_{\alpha}$ at the top. In practice, many components of $m^{(n)}_{\alpha}$ will be zero, as the embedding stores only features observed during the replay of word $\alpha$.
To compress this representation, we apply a sequence of SVD decompositions layer by layer. First, we perform an SVD of the capped tensor:
(32) $m^{(n)}_{\alpha,\mu_{n}}=\sum_{\beta_{n}}\hat{m}^{(n)}_{\alpha,\beta_{n}}\,V_{n}^{\beta_{n},\mu_{n}},$
where $V_{n}$ defines a word-specific low-rank subspace over $\mu_{n}$, and $\hat{m}^{(n)}_{\alpha,\beta_{n}}$ is the compressed top cap.
Next, we contract $V_{n}$ with $P_{n}$:
(33) $\tilde{P}_{n}^{\beta_{n},\mu_{n-1},j_{n}}=\sum_{\mu_{n}}V_{n}^{\beta_{n},\mu_{n}}P_{n}^{\mu_{n},\mu_{n-1},j_{n}}.$
This new projector has reduced dimension, containing only $n$-grams relevant to $a_{\alpha}$. We then reshape this tensor into a matrix over index pairs $(\mu_{n-1},j_{n})$ and perform another SVD: $\tilde{P}_{n}^{\beta_{n},(\mu_{n-1},j_{n})}=\sum_{\omega}U_{n}^{\beta_{n},\omega}\,D_{n}^{\omega}\,V_{n}^{\omega,(\mu_{n-1},j_{n})}$.
We now proceed to compress the next layer. We contract the decomposed $V$ from above with the next projector:
(34) $T_{\omega,j_{n},\mu_{n-2},j_{n-1}}=\sum_{\mu_{n-1}}V_{n}^{\omega,\mu_{n-1},j_{n}}\,P_{n-1}^{\mu_{n-1},\mu_{n-2},j_{n-1}}.$
Merging the left and right index pairs, we perform another decomposition: $T_{(\omega j_{n}),(\mu_{n-2}j_{n-1})}=\sum_{\beta_{n-1}}U^{\prime}_{(\omega j_{n}),\beta_{n-1}}D^{\prime}_{\beta_{n-1}}V^{\prime}_{\beta_{n-1},(\mu_{n-2},j_{n-1})}$. We now define the fully compressed projector:
(35) $\hat{P}_{n}^{\beta_{n},\beta_{n-1},j_{n}}$ $=\sum_{\omega}U_{n}^{\beta_{n},\omega}\,U^{\prime}_{\omega,j_{n},\beta_{n-1}}\,D^{\prime}_{\beta_{n-1}},$
(36) $\tilde{P}_{n-1}^{\beta_{n-1},\mu_{n-2},j_{n-1}}$ $=V^{\prime}_{\beta_{n-1},\mu_{n-2},j_{n-1}}.$
We repeat this contraction-decomposition process recursively, defining new compressed projection maps $\hat{P}_{n}$ and reduced intermediate indices $\beta_{n-1},\beta_{n-2},\dots$ at each step.
The final expression for the replayed memory becomes:
(37) $\sum_{\beta_{n},\dots,\beta_{2}}\hat{m}^{(n)}_{\alpha,\beta_{n}}\,\hat{P}_{n}^{\beta_{n},\beta_{n-1},j_{n}}\cdots\hat{P}_{2}^{\beta_{2},j_{1},j_{2}}\,v_{j_{1}}v_{j_{2}}\cdots v_{j_{n}}.$
This compressed chain retains only the information needed to reconstruct the replay trajectory of word $\alpha$, eliminating all unrelated feature branches. The compression is exact if no finite SVD eigenvalues are discarded. Thus this naturally generalizes to approximate compressions via truncating irrelevant contributions to the SVD spectra. We leave this latter point for future discovery.
Benefits.
This word-specific decomposition achieves two goals:
  • Disentanglement: Removes unrelated replay trajectories (an exact compression).
  • Compression: Optimizes the dimension of projection maps; & can regulate embedding dimension (approx).
5.5. Super-Hierarchical Key-Value Memory
The hierarchical key-value memory mechanism naturally extends to higher levels of abstraction. Just as short-term memory tokens $\tilde{v}^{(n)}$ are bound into long-term word embeddings $a_{\alpha}$ via replay, the embeddings themselves may form higher-order structures, such as phrases, idioms, or conceptual clusters. We introduce a second memory hierarchy built on projected embedding neurons $\tilde{a}^{(m)}$, with feature maps $\mathcal{P}_{m}$, correlation weights $G^{(m)}$, and memory tensors $M^{(m)}$. This forms a super-hierarchical memory, enabling compositional inference over learned embedding structures.
To support compositional memory beyond single-word representations, we extend the Hamiltonian to include higher-level interactions among embedding neurons:
$H$ $=\sum_{n}\sum_{\mu_{n-1},k}g^{(n)}_{\mu_{n-1},k}\tilde{v}^{(n-1)}_{\mu_{n-1}}v_{k}+\sum_{\alpha}a_{\alpha}\left(\sum_{n}\sum_{\mu_{n}}m^{(n)}_{\alpha,\mu_{n}}\tilde{v}^{(n)}_{\mu_{n}}\right)$
(38) $\quad+\sum_{m}\sum_{\alpha_{m-1},\alpha}G^{(m)}_{\alpha_{m-1},\alpha}\tilde{a}^{(m-1)}_{\alpha_{m-1}}a_{\alpha}+\sum_{\xi}b_{\xi}\left(\sum_{m}\sum_{\alpha_{m}}M^{(m)}_{\xi,\alpha_{m}}\tilde{a}^{(m)}_{\alpha_{m}}\right),$
where $a_{\alpha}$ are embedding neurons and $b_{\xi}$ are super-embedding neurons. Feature projections $\tilde{v}^{(n)}$ and $\tilde{a}^{(m)}$ are defined via retokenization hierarchies $P_{n}$ and $\mathcal{P}_{m}$ respectively.
Update Logic:
We compute gradients
(39) $\Delta v_{k}$ $=\frac{\partial H}{\partial v_{k}},\quad\Delta a_{\alpha}=\frac{\partial H}{\partial a_{\alpha}},\quad\Delta b_{\xi}=\frac{\partial H}{\partial b_{\xi}},$
and apply bounded updates: $Q\leftarrow f_{Q}(\Delta Q)$ for $Q\in\{v_{k},\ a_{\alpha},\ b_{\xi}\}$. We treat $b_{\xi}$ $a_{\alpha}$ as being faster than $v_{k}$ (i.e hierarchy of equilibriation times $\tau_{b}\ll\tau_{a}\ll\tau_{v}$ (Krotov and Hopfield, 2021; Krotov, 2021)). Thus we solve for $b_{\xi}$ and $a_{\alpha}$ first, then substitute into $\Delta v_{k}$.
Inference Equation:
Letting $g^{(n)}=G^{(m)}=0$, we derive the inference rule:
(40) $\Delta v_{k}=\sum_{\alpha}f_{\alpha}\left(\sum_{n^{\prime}}\sum_{\mu_{n^{\prime}}^{\prime}}m^{(n^{\prime})}_{\alpha,\mu_{n^{\prime}}^{\prime}}\tilde{v}^{(n^{\prime})}_{\mu_{n^{\prime}}^{\prime}}+\sum_{\xi}f_{\xi}\left(\sum_{m^{\prime}}\sum_{\alpha_{m^{\prime}}^{\prime}}M^{(m^{\prime})}_{\xi,\alpha_{m^{\prime}}^{\prime}}\tilde{a}^{(m^{\prime})}_{\alpha_{m^{\prime}}^{\prime}}\right)\right.\\ \left.\times\sum_{\alpha_{m-1}}M^{(m)}_{\xi,\alpha_{m-1},\alpha}\tilde{a}^{(m-1)}_{\alpha_{m-1}}\right)\sum_{n}\sum_{\mu_{n-1}}m^{(n)}_{\alpha,\mu_{n-1},k}\tilde{v}^{(n-1)}_{\mu_{n-1}}.$
This defines a super-hierarchical key-value memory: the context selects a super-embedding $b_{\xi}$, which activates relevant embeddings $a_{\alpha}$, which in turn retrieve fine-grained retokenized features for predicting $v_{k}$.
5.5.1. Plasticity, Emergence, and the role of Space & Time
The emergent key-value memories discussed here are a consequence of the plastic nature of Hebbian learning, combined with the hierarchical tokenizing inference. These new mechanisms are not put in by hand – there is no “key-value memory layer” added to the system – and no retraining of the entire model is necessary. Instead, introducing a single new neuron to the system is sufficient in order for them to arise, demonstrating that they are truly emergent structures. (Nor is explicit data necessary – see Discussion.)
This type of plasticity is not present in models which learn via optimizing a global objective (back propagation). While such models are typically modular, their architecture is fixed beforehand. Though the black box paradigm permits a type of data-driven “emergence”, it is not the type of structural emergence described here, but instead reflects a discovery of internal representations allowed within the fixed architecture. The attention mechanism of an LLM is assumed beforehand (Vaswani et al., 2017), not a learned structured. This differs from Conway’s Game of Life, which can “learn” a Turing machine – not from recognizing patterns in data nor minimizing an objective, but as an accident due to the local dynamics & initial conditions (Rendell, 2011, 2014).
The hierarchical key-value memories described here can be understood as arising due to the interplay of space & time. The spatial aspect is reflected in how the key-value relationship preserves the morphological structure, such that the value part encodes how to complete a word from a triggered key. Thus no positional embeddings are necessary (or helpful). This spatial awareness is learned because retokenization makes replay boundary aware; so that the added embedding neuron correlates with the features corresponding to smoothly correlated regions of string. By having the neuron activated throughout this cycle, it learns this smooth feature set as if all its elements fired simultaneously during a single Hebbian update step. At face value this process is “precisely timed”, but thanks to the spatial awareness of retokenization, doing this precise timing is trivial.
It’s useful to think of replay as a minimally structured game (or a local dynamic coding scheme) – grab a new neuron, poke short-term memory to randomly replay its correlations, then repeat the process again with a new neuron once replay terminates. It works because the local event learning makes continued learning & emergence incidental. The assumed structure of replay is chosen to exploit this “accidental learning”. Likely other (more efficient less random) structured games exists which exploit this accident.
Even following our minimal replay blueprint, the number of possible different emergent structures grows combinatorially with the number of learned long-term embeddings. As more embedding neurons populate the long-term memory ($V_{\Omega}:=(v_{k},a_{\alpha},b_{\xi},\cdots)_{\Omega}$), the model can learn correlations between these embeddings ($V_{\Omega}V_{\Omega^{\prime}}$ for subset of $\Omega$), which can then form entirely novel compositional structures (beyond hierarchical & super-hierarchical). These newly learned structures, in turn, enrich the token representations available to short-term memory, enabling further complexity. Classifying all possible emergent structures, given the local dynamics, is akin to exactly solving a physics problem. In general this is not always possible, but nevertheless opens a potentially fruitful direction for future research.
Figure 3. Hierarchies of hierarchies: STM’s are n-gram models which build features ($\tilde{v}$,$\tilde{a}$,$\tilde{b}$,…) from relevant correlations ($\tilde{v}v$,$\tilde{a}a$,…). These can be replayed into symbolic key-value memory tokens of increasing complexity LTM ($a$,$b$,…). We show the emergent inference classes learned at each step. In addition to the direct hierarchies, LTM tokens of mixed class can be correlated.
6. Toy Example: Finger Counting
This example illustrates how replay and hierarchical memory enable the model to bind structured inputs (e.g. finger gestures) to symbolic tokens (e.g. number words).
6.1. Modalities and Embeddings
We assume two input modalities:
  • Letter modality: Basis tokens $v_{k}$ for alphabet letters (e.g., $v_{\bf o},v_{\bf n},v_{\bf e}$).
  • Finger modality: Scalars $u\in\{0,1\}$ indicating raised fingers.
Each embedding $a_{\alpha}$ is associated with hierarchical memory:
$m_{\alpha}=\sum_{n}\sum_{\mu_{n}}m^{(n)}_{\alpha,\mu_{n}}\tilde{v}^{(n)}_{\mu_{n}}$
For example, the word “one” is stored as:
$m_{1}$ $=$ $m_{1,\bf o}^{(1)}v_{\bf o}+m_{1,\bf n}^{(1)}v_{\bf n}+m_{1,\bf e}^{(1)}v_{\bf e}$
$+$ $m_{1,\bf on}^{(2)}\tilde{v}_{\bf on}^{(2)}+m_{1,\bf ne}^{(2)}\tilde{v}_{\bf ne}^{(2)}+m_{1,\bf one}^{(3)}\tilde{v}_{\bf one}^{(3)},$
and similarly for “two” and “three”.
Finger gestures are retokenized as:
$\tilde{u}^{(n)}=\prod_{x=1}^{n}u_{x},$
with $\tilde{u}^{(n)}=1$ if $n$ fingers are raised.
6.2. Binding Fingers to Numbers
We define the Hamiltonian:
(41) $H$ $=\sum_{\alpha}a_{\alpha}\left(\sum_{n}\sum_{\mu_{n}}m^{(n)}_{\alpha,\mu_{n}}\,\tilde{v}^{(n)}_{\mu_{n}}\right)$
(42) $\quad+G_{1}^{(1)}\,a_{1}\,u+G_{2}^{(2)}\,a_{2}\,\tilde{u}^{(2)}+G_{3}^{(3)}\,a_{3}\,\tilde{u}^{(3)}.$
where the $G$ couplings are learned via Hebbian updates, e.g. showing two and two raised fingers. (Technically, replay will bind $\alpha>n$ embeddings to $\tilde{u}^{(n)}$, such as $a_{3}$ being bound to $u$, $\tilde{u}^{(2)}$, and $\tilde{u}^{(3)}$. But this interference is mitigated by biasing $G_{\alpha}^{(n)}\rightarrow\beta^{n}G_{\alpha}^{(n)}$ with $\beta>1$, and choosing $f_{\alpha}$ to be normalization.)
During inference we raise 2 fingers ($\tilde{u}^{(n)}=\sum_{n^{\prime}\leq 2}\delta_{n^{\prime},n}$),
$\Delta a_{\alpha}=G_{\alpha}^{(n)}\tilde{u}^{(n)}\quad\Rightarrow\quad a_{\alpha}\mapsto\text{``two''}\Rightarrow v_{k}=t,w,o.$
That is, a finger count activates the corresponding embedding, which then drives autoregressive symbol generation following the key-value relationship encoded in Eqn 41.
7. hierarchical feature mixing & random language
We begin by analyzing the structure of a randomly sampled vocabulary, consisting of 100 personal names (in English or Spanish) (dat, aset). For this set, we plot the $d_{n}$ distribution, where $d_{n}$ counts the number of unique $n$-grams observed at each scale. The resulting $d_{n}$ distribution follows an approximately log-normal form, with the largest $n$ determined by the longest name in the sample.
To probe the internal structure of the learned hierarchy, we perform a game of cut-and-regrow (Fig 5). Specifically, we impose a cutoff $n_{\text{cut}}$ and delete all $n$-gram features with $n\geq n_{\text{cut}}$, then allow the hierarchy to regrow under the smoothness constraint. This process reveals how the underlying directed acyclic graph (DAG) of features governs regrowth: surviving features at small $n$ serve as anchoring nodes, from which higher-order structures emerge. Memory is locally organized, such that the shortest word-length scales dominate.
When cutting at $n_{\text{cut}}\gg n_{\text{peak}}$, where $n_{\text{peak}}$ is the location of the original $d_{n}$ peak, regrowth closely reproduces the original structure, introducing minimal (or no) novelty. This is because the many lower layers of the hierarchy provide strong constraints on the space of learnable regrowths.
In contrast, cutting closer to $n_{\text{peak}}$ loosens the scale-dependent constraints, allowing novel $n$-grams to form. Nevertheless, even here, the surviving short-scale DAG structure continues to guide regrowth, ensuring that generated words remain smooth and morphologically coherent.
We find that though the original notion of smoothness is perfectly preserved below the cut, cutting away more of the $d_{n}$ peak causes the hierarchy to begin losing its recognizable English/Spanish nature. This highlights the importance of the $n_{\text{peak}}$ grams to language structure [^5] For $n\lesssim n_{\text{peak}}$, $d_{n}$ grows approximately exponentially, with learned constraints ($g^{(n)}=0$) pruning branches and thereby reducing the effective growth rate. For $n>n_{\text{peak}}$, the smoothness condition takes over: new tokens are permitted only if all subgrams are valid, which recursively restricts composition and causes a collapse in $d_{n}$. This collapses guarantees finite word length (Eugenio, 2025). In formal language terms, this might be described as a consequence of the regular-language-like limitations of stored strings..
Finally, if we cut at $n_{\text{cut}}=2$, we discard all memory of the original names. This is the data-free regime, where growing a hierarchy generates what we refer to as a “random language” (Eugenio, 2025). Nevertheless, growing hierarchies under the smoothness constraint still yields finite word-lengths and an approximately log-normal $d_{n}$ distribution (HERDAN, 1958; WILLIAMS, 1940; Limpert et al., 2001). This shows that bounded, hierarchical morphology is an intrinsic feature of the model itself, independent of any training data, and that language structure arises naturally from smooth hierarchical learning.
Further, this behavior illustrates a key property of the model: Short-term memory structures $g^{(n)}$ behave like a dynamic scratch board, where hierarchical feature mixing enables creative recombination under smoothness preservation. The model’s DAG structure naturally channels this process, enforcing compositionality while permitting exploration of new morphological forms.
7.1. Synaptic Zipf-laws
Random languages generated by our model exhibit rank-ordered $n$-gram frequency distributions that follow a scale-invariant Zipf-like law (Zipf, 1949; Gan et al., 2009), a signature of morphological structure (Eugenio, 2025). These spectra provide insight into the statistical relevance of subword patterns, ranked from most to least frequent, and reveal how this structure scales with increasing pattern length.
Note we can interpret these ranked frequencies as defining an effective memory spectra $g^{(n)}_{\text{eff}}$, where Hebbian learning imprints frequency counts from local events $\mathcal{E}_{n}$ directly into memory:
$g^{(n)}_{\text{eff}}=\sum_{\mathcal{E}_{n}^{\small\text{vocab}}}\tilde{v}v.$
In the absence of forgetting (Eugenio, 2025), this spectrum reflects all observed correlations of length $n$.
Fig 4 shows the effective model spectrum corresponding to a model whose memory graph was grown randomly under the hierarchical DAG constraint. As such, the resulting distribution reflects only the inductive bias of the memory structure itself.
At short lengths ($n<n_{\text{peak}}$), constraints are minimal, and the distribution appears structureless: a Zipf-like power-law head followed by a flat tail of irrelevant $n$-grams, similar to those found in uniform random strings (Eugenio, 2025; Gan et al., 2009). However, as $n$ increases, the DAG constraint increasingly shapes memory formation. Beyond the peak, the tail vanishes and the power-law scaling persists, signaling the emergence of a dominant compositional structure at longer lengths.
Figure 4. The frequency an $n$-gram (length $n$) appears in the vocabulary, rank-ordered from most to least frequent. The vocabulary is for the random language bjfyi (in supplement). Equivalently, this plot can be understood as representing an effective model spectra ($g^{(n)}_{\text{eff}}$ per $n$). Scale-invariance emerges beyond $n_{\text{peak}}=3$ due to the compositional structure.
8. Discussion
8.1. Where did the data come from?
Neither punctuation nor spacing are necessary elements of writing:
[…]
This property is not special to this example, but is a general quality of human language patterns. Historically, human writing existed without specialized tokens indicating boundary. This is perhaps best illustrated by alphabets, whose fundamental tokens are meaningless beyond their sound associations. And yet such texts could (and can) still be read. This is only possible because human language is inherently tokenizable. There exists a statistical distinction between word-forming and non-word-forming patterns. The later introduction of spaces, punctuation, and capitalization served to mark these boundaries, but they are not themselves fundamental to the structure of language. They are external annotations for a structure already present in the stream of phonetic or orthographic symbols.
The existence of tokenizable structure in language raises a deeper question: why should such structure exist at all? Conventional language models don’t address this distinction. All tokens—whether letters, spaces, punctuation, or stop tokens—are treated as elements of a single, undifferentiated vocabulary. These models assume the data has already been segmented, and that its structure reflects a meaningful set of distinctions. But this sidesteps the question of origin. The model is not responsible for discovering the structure of language, only reproducing it. By assuming the existence of a token vocabulary and training on massive amounts of pre-tokenized data, the model inherits the statistical regularities of language without explaining them. It becomes a powerful recognizer, but not a theory of how language—and its specific structure—emerges at all.
To study how such structure can emerge, we present our model with the 26 lowercase alphabetic characters as its only input features. This is not a simplification, but the intended use of the model. Learning in our framework proceeds by growing a memory structure dictated by local Hebbian updates constrained to follow a hierarchical DAG. Information is retained only when the correlations in the data align with what the local memory can represent. Presenting the model with one word at a time, or separating words by zero vectors rather than an actual token, allows it to build a memory that respects the data structure. If a space character were included as an actual token, it would be treated like any other feature, leading the model to learn correlations involving it. This would break locality, and the model would no longer distinguish between word-forming and non-word-forming sequences. Space would be learned as part of words, not as a boundary, and the accepted strings of the model would no longer correspond to linguistic units. The model would simply learn a different interpretation of the data—one that reflects statistical artifacts of the tokenization, not the underlying structure of language.
This is not a deficiency, but a reflection of how learning occurs in the model. Because it has no global objective, and no access to the meaning of individual symbols, the model aligns new information with structures that have already emerged from past correlations. When trained on unstructured noise—such as a uniform string of randomly chosen letters—the model still forms memory traces. It does so by capturing those accidental patterns in the noise that align with the allowed DAG growth. The result is a memory of accepted strings whose collective statistical properties closely match the morphological structure of natural language. In this sense, the model is capable of discovering natural linguistic patterns without data, because its internal memory structure mirrors the same constraints that shape human word formation (Eugenio, 2025). Including a space token misleads the model, not because it fails to recognize it, but because it has no way of knowing that such a symbol is meant to indicate a boundary, rather than a letter.
Thus it provides a scientific explanation for the universal tokenizability of human language, not just because it processes language data in a locally coordinated (and therefore biologically plausible) way, but because it can tell us why the data exists in the first place.
8.2. Scaling Up
Scaling up the model introduces unique challenges. Naively treating the retokenizer as a conventional tokenizer—such as byte-pair encoding (BPE) (Gage, 1994) or other subword segmentation (Schuster and Nakajima, 2012)—misses the point entirely. The retokenizer is not a preprocessing step, but a dynamic process. It functions as a compositional short-term memory system, one that stores and retrieves accepted strings by activating memory traces across the hierarchy. Discarding this hierarchical structure and collapsing all learned $n$-grams into a flat vocabulary would scale poorly and break the model’s inductive biases. The purpose of the retokenizer is not to produce tokens for input to another model, but to generate new feature bases through replay, enabling creative learning by aligning input structure with emergent memory.
While subword segmentation methods such as byte-pair encoding (BPE) (Gage, 1994) and unigram language modeling (Kudo, 2018) have become standard in NLP, they rely on statistical preprocessing over large corpora. A variety of recent approaches attempt to learn token boundaries from data, rather than relying on fixed preprocessing rules. These include segmentation-aware training using multiple subword candidates (Kudo, 2018), latent segment models that infer tokenization jointly with downstream tasks (Anonymous, 2024), and byte-level systems that learn token patches directly from raw input (Pagnoni et al., 2024; Schmidt et al., 2025). Probabilistic methods such as Bayesian word segmentation (Goldwater et al., 2009) and morphology induction algorithms like Morfessor (Creutz et al., 2006) also aim to discover subword units in an unsupervised way. Yet all of these approaches depend fundamentally on the assumption that the data contains language. That is, the model must learn to tokenize from patterns already present in the input. If no such structure exists—such as when trained on noise—these models fail to recover meaningful segmentations, because there is no statistical signal to exploit.
In contrast, the framework presented here does not learn token boundaries from data but instead defines them as a side-effect of emergent memory dynamics. It does this by forming accepted sequences through local Hebbian dynamics constrained by a hierarchical DAG. Tokenization is not assumed, but emerges as a side effect of correlation alignment. No loss is minimized, no boundaries are enforced, and no tokens are treated as intrinsically meaningful (e.g stop or space tokens). This allows the model to exhibit tokenizable structure even in the absence of language data—a capability absent in systems trained on the assumption that language structure is already embedded in the dataset.
Going beyond the limitations of short-term memory thus requires solutions which don’t ruin the emergence. Replay provides such a mechanism, as its minimal structure exploits the local dynamics to grow the model in a useful way. It suggests a framework where learned tokens are correlated in short-term memory structures, whose correlations are replayed in order to contribute yet more complex tokens into long-term memory. This problem is complicated by the lack of static architecture. This makes “scaling up” a game of assuming and exploring allowed structures, including which neurons are allowed to correlate in short-term memories, and the order and manner of replay.
Compression likely plays a critical role in scaling emergent structure. In this work, we focused on the simplest and most analytically tractable setting: binding subword patterns into word embeddings. Short-term memory need not store every $n$-gram, and the few-word examples we present are representative of general behavior. We apply an exact matrix product state (MPS) compression—equivalent to a tensor train decomposition—without truncating singular values, yielding fully disentangled feature maps and one embedding token per word. The resulting long-term memory functions as an attention-like, one-to-one key-value map over these discrete embeddings.
This exact decomposition demonstrates how structured embeddings can emerge from local memory dynamics. But the same MPS protocol naturally supports approximate compression by truncating small singular values, allowing generalization across statistically similar tokens. While we reserve empirical study of approximate compression for future work, its success in many-body physics (Gourianov et al., 2022; Schollwöck, 2011; Parker et al., 2020; Stoudenmire, 2022; Gourianov et al., 2025; Gourianov, 2022), and its growing adoption in machine learning under the name tensor train (Xu et al., 2024; Su et al., 2024; Novikov et al., 2015; Hrinchuk et al., 2020), suggest that it may be essential for scaling this framework beyond the low-data regime.
8.3. Static vs Dynamic Learning
Lastly, while the incompatibility between emergent learning and modern NLP frameworks poses challenges for scaling and experimentation, it raises a more fundamental question: Are we trying to build models of cognition – capable of continual learning, inter-generational memory, and emergent language capacity – or are we refining static pattern recognizers? These are separate goals rooted in different paradigms of intelligence.
Current language models operate in a globally optimized regime: they treat learning as a statistical estimation problem, where the architecture is fixed and the learning dynamics are frozen into a high-dimensional variational ansatz. In this sense, such models are effectively thermalized – time plays no intrinsic role in learning.
Consider that “thinking” in static models means test-time compute – but in human learning, thinking is a dynamic learning process, where observed events are reprocessed. Though advances in reinforcement learning demonstrate how a pre-trained LLM can learn by reasoning to itself (Zhao et al., 2025b, a), such zero-data learning is only possible because of the pretraining, and therefore not integral to how the model learns.
This contrasts with local event-driven learners, which require thinking for the model to grow beyond its initial locally-coordinated limitations. Locally-learned events are little more than simple replayable memories stored in an uncoordinated way. Memory retrieval is therefore essential to continued learning, not just because it’s useful to do so, but because it can do little else without further coordination.
8.4. Accident of Nature
Beyond the desire to advance technology, there is also the scientific question of human cognition. Global gradient descent presupposes (1) that a global notion of correctness exists (e.g a loss function), and (2) that neurons coordinate their updates to minimize it (via teaching signals (Maoz et al., 2020)). Such global coordination is not only incompatible with the microscopic behavior of neurons, but may not be a description of any naturally occurring microscopic system. The assumption of global coordination violates a basic assumption of physics: that nature is an unconscious collection of accidents.
Attempts to justify LLMs as biologically plausible ignore this disconnect: a black-box model that passively reflects its data distribution cannot offer a generative theory of language, only an empirical fit to it. This makes drawing scientific conclusions impossible, thus leaving fundamental questions unanswered: Why is speech segmented into words? What enforces the universality of word-forming patterns? How does linguistic structure emerge in the first place, and what makes it stable across generations? Even at the low-level explored in this work, our framework addresses these questions.
9. conclusion
We have introduced a fully local, objective-free framework for hierarchical language learning, where compositional structure emerges from unsupervised retokenization and replay. Rather than fitting to data, the model constructs an internal, interpretable hierarchy of features, enabling smooth, bounded morphological structures to arise even in the absence of training data.
This framework provides natural mechanisms for scaling up. Through smooth hierarchical composition and replay-based memory formation, the model can build long-range correlations, compress complex structures into efficient key-value memories, and recursively retokenize features across multiple scales. In principle, this enables the emergence of increasingly complex symbolic systems from simple local interactions.
However, scaling up without a global objective is not straightforward. Without loss functions to optimize, the growth of complexity depends on the evolutionary dynamics of replay, memory, and hierarchical interactions—analogous to how structure emerges in systems like Conway’s Game of Life (Rendell, 2011, 2014; Codd, 1968) or Lenia (Chan, 2019; Plantec et al., 2023; Faldor and Cully, 2024; Chan, 2020). Progress toward richer, longer-range structure may require evolutionary strategies: introducing multiple interacting sub-networks, varying timescales of memory and replay, and selecting architectures that favor the extension of correlations over time.
A crucial distinction of this approach is that the model does not fit data – it interprets data. The system is fundamentally biased toward smooth, hierarchical DAG-like representations (Solan et al., 2005; Beavers, 2010; Beavers and Harrison, [n. d.]), which govern its construction of internal structure. This intrinsic bias explains its ability to extract natural language morphology from noise, and suggests a fundamentally different paradigm of learning—one closer to human cognition, where models continually reinterpret the world rather than merely optimizing against a data distribution. This guarantees diversity in learning, where learned errors are reinterpreted in the context of intrinsic model structure, leading to language change instead of model collapse.
Ultimately, the broader goal is to build models that can develop an entire language from scratch, without preexisting data, by autonomously constructing a hierarchy of symbols, grammar, and meaning. By allowing multiple modalities of hierarchical learning—for example, connecting emergent symbol sequences with structured perceptual patterns like the finger-counting toy example—meaning can arise at the intersection of different feature sets tied together through replay and embedding. In this way, not only words but also concepts and their cross-modal relationships could emerge spontaneously, answering the question: “How did we go from a universe without language to a universe with language?”
Figure 5. Number of unique $n$-grams in a 100-name dataset (English/Spanish mix) (dat, aset) shown as backdrop. Shaded regions mark features removed by hierarchical cuts; the solid line fits a log-normal distribution. Example words from post-cut regrowth are shown.
Code Availability
Acknowledgements
We thank Larry Moss, Ricky Simanjuntak, & Tony Beaver for comments/discussion.